OpenStudy (anonymous):

In the 1860's, on a farm along the Mississippi River, a girl was presented with three gifts: a pig, a goose, and a duck. The duck had cost 20 cents. The pig and the goose together cost twice as much as the duck. The goose and the duck together cost three times as much as the pig. How much did the goose cost alone?

OpenStudy (anonymous):

The duck had cost 20 cents d = 20 The pig and the goose together cost twice as much as the duck p + g = 40 The goose and the duck together cost three times as much as the pig g + 20 = 3p Use substitution with those last 2 equations P + g = 40 ======> p = 40 - g Plug that into the last equation g + 20 = 3(40 - g) Solve that for g

OpenStudy (anonymous):

thanks,i'll try this

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

is the answer 10 cents?because the pig must be 30 cents,so30+10=40

OpenStudy (anonymous):

Refer to @BangkokGarrett's solution's last equation: g + 20 = 3(40 - g) and solve it for g.

OpenStudy (anonymous):

Is it g=30 i don't get it mabee because i am in 6ht grade.

OpenStudy (anonymous):

g-20=3(40-g) g-20+20=120-g+20 I do not think this is correct